OpenStudy (anonymous):

Is 0 a solution of -2 < x < 5? A. yes B. no

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Is the following statement true?\[ -2<0<5 \]Try drawing it on the number line.

OpenStudy (anonymous):

So it is A

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thanks :D

OpenStudy (anonymous):

0<5 0>-2

OpenStudy (ajprincess):

The solutions are -1, 0, `1, 2, 3 and 4

OpenStudy (anonymous):

wow you are thorough Princess :)

OpenStudy (anonymous):

@aj that's only true if we're working in \(\mathbb N\)

OpenStudy (anonymous):

If we're in \(\mathbb Q\) or \(\mathbb R\) there are more solutions than that.

OpenStudy (ajprincess):

ya. I agree vth u @nbouscal

OpenStudy (anonymous):

(-2,5)